# How to Create a Flat Ten or Eleven Dimensional Space-time in the Interior of an Asymptotically Flat Four Dimensional String Theory

### v1

@@ Abstract
By taking large mass and charge limit of a black hole in string theory we can create arbitrarily large regions where the space-time is approximately flat, but the moduli fields take values different from their asymptotic values. In this paper we describe a special case of this where black hole solutions in a four dimensional string theory, in the large mass and charge limit, can have an arbitrarily large region outside the horizon where a local observer will experience string theory in flat ten dimensional space-time. The curvature and other field strengths remain small everywhere between the asymptotic four dimensional observer and the ten dimensional region. By going to a different region of space, we can also get a large region where a local observer experiences M-theory in flat eleven dimensional space-time.
@@ 1 Introduction
One example of this was discussed in [1] where it was shown how a large black hole in ten dimensional type IIA string theory carrying D0-brane charge allows us to create arbitrarily large, almost flat regions where the string coupling remains almost constant at any value between its asymptotic value and infinity. A large string coupling would correspond to M-theory in an almost flat eleven dimensional space-time[3, 4]. In this paper we describe another example where we begin with type IIA string theory compactified on $T^{6}$ and produce a background that contains a large, locally flat region where the size of the internal $T^{6}$ can be made arbitrarily large, keeping the string coupling fixed. Therefore a local observer in this region will experience type IIA string theory in ten dimensional flat space-time. There is another region in the same background where a local observer experiences M-theory in eleven dimensional flat space-time. The only price we pay is that larger the size of this ten or eleven dimensional region, the larger is the mass of the black hole needed to create such a configuration.
The rest of the paper is organized as follows. In section 2 we describe a background where we have a locally flat region with large internal $T^{6}$, but the string coupling in that region is also large. This is not quite the configuration we want since we want the string coupling to remain small or finite. We remedy this problem in section 3 by placing the black hole solution of section 2 in the background of a second black hole. There we show that it is not only possible to identify a large region where the local physics is described by type II string theory in ten dimensional flat space-time, but it is also possible to identify another large region where the local physics is described by M-theory in eleven dimensional flat space-time. We end in section 4 with some comments.
@@ 4 Discussions
Finally we note that while the asymptotic geometry is taken to be flat space-time (times an internal manifold) the geometries that we can construct in the interior are not limited to flat space-time. For example is ten dimensional type IIB string theory we can take a stack of N spherical D3-branes so that it has finite energy, but take the radius of the sphere to be very large so that the relaxation time is large. In that case the near horizon geometry of a local patch of the D3-branes will be $AdS_{5}\times S^{5}$ with five form flux[9] and we can study properties of this background by sending observers close to the horizon. [^5] If we scale the 4-form field strength according to (2.5) then the three-brane charge will grow as $\lambda^{4}$[1] and we shall get a locally flat space-time. To get $AdS_{5}\times S^{5}$ geometry we need to keep the three brane charge fixed. Of course if we try to approach infinitely close to the horizon any experiment will take infinite amount of time from the asymptotic observer’s point of view due to the red shift, but once we fix the accuracy with which we want to do the experiment and estimate the time that will be needed from the asymptotic observer’s point of view to perform such an experiment, we can take the size of the 3-brane system to be sufficiently large so that it does not evolve appreciably during the period of the experiment.

### v2

@@ Abstract
By taking large mass and charge limit of a black hole in string theory we can create arbitrarily large regions where the space-time is approximately flat, but the moduli fields take values different from their asymptotic values. In this paper we describe a special case of this where black hole solutions in a four dimensional string theory, in the large mass and charge limit, can have an arbitrarily large region outside the horizon where a local observer will experience type IIA string theory in flat ten dimensional space-time. The curvature and other field strengths remain small everywhere between the asymptotic four dimensional observer and the ten dimensional region. By going to a different region of space, we can also get a large region where a local observer experiences M-theory in flat eleven dimensional space-time. By taking another solution in the same theory, one can create an arbitrarily large region where a local observer will experience type IIB string theory in flat ten dimensional space-time.
@@ 1 Introduction
One example of this was discussed in [1] where it was shown how a large black hole in ten dimensional type IIA string theory carrying D0-brane charge allows us to create arbitrarily large, almost flat regions where the string coupling remains almost constant at any value between its asymptotic value and infinity. A large string coupling would correspond to M-theory in an almost flat eleven dimensional space-time[3, 4]. In this paper we describe another example where we begin with type IIA string theory compactified on $T^{6}$ and produce a background that contains a large, locally flat region where the size of the internal $T^{6}$ can be made arbitrarily large, keeping the string coupling fixed. Therefore a local observer in this region will experience type IIA string theory in ten dimensional flat space-time. There is another region in the same background where a local observer experiences M-theory in eleven dimensional flat space-time. The only price we pay is that larger the size of this ten or eleven dimensional region, the larger is the mass of the black hole needed to create such a configuration. Finally, by taking a different background in the same theory, one can construct an arbitrarily large region where a local observer experiences type IIB string theory in ten dimensional flat space-time.
The rest of the paper is organized as follows. In section 2 we describe a background where we have a locally flat region with large internal $T^{6}$, but the string coupling in that region is also large. This is not quite the configuration we want since we want the string coupling to remain small or finite. We remedy this problem in section 3 by placing the black hole solution of section 2 in the background of a second black hole. There we show that it is not only possible to identify a large region where the local physics is described by type IIA string theory in ten dimensional flat space-time, but it is also possible to identify another large region where the local physics is described by M-theory in eleven dimensional flat space-time. In section 4 we show, how by taking a different background in the same theory, we can produce a large region where a local observer experiences type IIB string theory in ten dimensional flat space-time with moderate or small string coupling. We end in section 5 with some comments.
@@ 3 Black hole with winding and momentum charge
4 Black hole with purely winding charge
In this section we shall describe how to construct a configuration that contains an arbitrarily large region where the local observer experiences type IIB string theory in flat ten dimensional space-time. For this we need to make use of a third black hole solution in type IIA string theory on $T^{6}$ that carries only winding charge along one of the circles. We shall take this to be the ninth circle. This solution can again be read out from the general solution given in [6]. [^5] The solution corresponds to setting $\alpha=\beta$ in [6]. The radius $R_{9}$ of the ninth circle appears inside the matrix $M$ as a $2\times 2$ bloc matrix of the form ${1\over 2}\pmatrix{R_{9}^{2}+R_{9}^{-2}&R_{9}^{2}-R_{9}^{-2}\cr R_{9}^{2}-R_{9}^{-2}&R_{9}^{2}+R_{9}^{-2}}\,.$ ) . Using this the four dimensional dilaton $\phi_{4}$ and the radius $R_{9}$ of the ninth direction associated with the solution in [6] for $\alpha=\beta$ takes the form:
$e^{2\phi_{4}}=e^{2\widehat{\phi}_{0}}\,(1+2\widehat{m}\widehat{\rho}^{-1}\sinh^{2}\alpha)^{-1/2}\,,$ (4.1)
$R_{9}=(1+2\widehat{m}\widehat{\rho}^{-1}\sinh^{2}\alpha)^{-1/2}\,,$ (4.2)
where we have introduced new parameter $\widehat{m}$ and new radial variable $\widehat{\rho}$ to distinguish them from the variables used in section 3. The other radii are kept fixed at 1. ${\widehat{\phi}_{0}}$ denotes the asymptotic value of $\phi_{4}$ for this solution. – this is an extra parameter that was set equal to zero in [6]. This is also the asymptotic value of the ten dimensional dilaton $\phi$ for this solution since $R_{9}$ approaches 1 in this limit and the other radii are also fixed at 1. On the other hand at a generic position labelled by $\widehat{\rho}$, the ten dimensional dilaton $\phi$ takes the form
$e^{2\phi}=e^{2\phi_{4}}\,R_{9}=e^{2\widehat{\phi}_{0}}\,(1+2\widehat{m}\widehat{\rho}^{-1}\sinh^{2}\alpha)^{-1}\,.$ (4.3)
Let us now put this solution in the background of the solution given in (3) at the position $\rho$, taking the size of (3) to be much larger than the size of this solution (i.e. we take $m,\rho>>\widehat{m},\widehat{\rho}>>1$). Then we can use (3.2) to write
$e^{-2\widehat{\phi}_{0}}=m\,\rho^{-1}\,\cosh\beta\,.$ (4.4)
Taking $\widehat{m}\widehat{\rho}^{-1}\sinh^{2}\alpha$ to be large, we get, from (4.3), (4.2):
$e^{2\phi}={1\over 2}\,m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\widehat{m}^{-1}\,\widehat{\rho}\,(\sinh\alpha)^{-2}\,,\qquad R_{9}={1\over\sqrt{2}}\,\widehat{m}^{-1/2}\,\widehat{\rho}^{1/2}\,(\sinh\alpha)^{-1}\,.$ (4.5)
We now put the first black hole, described in section 2, at the position $\widehat{\rho}$ by taking $m,\rho>>\hat{m},\hat{\rho}>>r_{\pm},a,b>>1$. This amounts to replacing $e^{2\phi_{0}}$ in (2.9) by $e^{2\phi}$ given in (4.5). Also, the radius of the ninth circle will now be multiplied by $R_{9}$ given in (4.5) since this becomes the asymptotic value of this radius from the perspective of the solution (2.9). This gives the following solution:
$\widetilde{R}=\left({r_{+}+b\over b+a}\right)^{1/4},\qquad\widetilde{R}_{9}=\left({r_{+}+b\over b+a}\right)^{1/4}{1\over\sqrt{2}}\,\widehat{m}^{-1/2}\,\widehat{\rho}^{1/2}\,(\sinh\alpha)^{-1},$
$e^{2\widetilde{\phi}}={1\over 2}\,m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\widehat{m}^{-1}\,\widehat{\rho}\,(\sinh\alpha)^{-2}\,\left({r_{+}+b\over b+a}\right)^{3/2}\,.$ (4.6)
Note that if we ignore the ${1\over 2}\,\widehat{m}^{-1}\,\widehat{\rho}\,(\sinh\alpha)^{-2}$ factor, then the solution reduces to (3.3).
The final step is to make a $R\to R^{-1}$ duality transformation on the ninth circle to map this to a type IIB string theory. Denoting the new variables by bars, we get
$\overline{R}=\widetilde{R}=\left({r_{+}+b\over b+a}\right)^{1/4},\qquad\overline{R}_{9}=\widetilde{R}_{9}^{-1}=\left({r_{+}+b\over b+a}\right)^{-1/4}{\sqrt{2}}\,\widehat{m}^{1/2}\,\widehat{\rho}^{-1/2}\,\sinh\alpha\,,$
$e^{2\overline{\phi}}=e^{2\widetilde{\phi}}\widetilde{R}_{9}^{-2}=m^{-1}\,\rho\,(\cosh\beta)^{-1}\,\left({r_{+}+b\over b+a}\right)\,.$ (4.7)
In order to get type IIB string theory with large compactification radii and weak or moderate coupling, we need $\overline{R},\overline{R}_{9}>>1$ and $e^{\overline{\phi}}\stackrel{{\scriptstyle<}}{{\sim}}1$. If we choose $\rho\sim m$ and $\widehat{\rho}\sim\widehat{m}$, then this requires:
$r_{+}>>a,b,\qquad\sinh\alpha>>\left({r_{+}+b\over b+a}\right)^{1/4},\qquad\cosh\beta\stackrel{{\scriptstyle>}}{{\sim}}\left({r_{+}+b\over b+a}\right)\,.$ (4.8)
The analog of (3.5) now takes the form:
$m>>\widehat{m}>>r_{+}>>>\sinh^{4}\alpha>>r_{+}/(b+a)\sim\cosh\beta>>1,\qquad\rho\sim m,\qquad\widehat{\rho}\sim\widehat{m}\,.$ (4.9)
Finally, we note that if we had used the type IIB description from the begining, then the first black hole would carry D1-brane charge along the ninth direction, the second black hole would carry equal amount of fundamental string winding and momentum charges along one of the circles and the third black hole would carry momentum along the ninth direction.
@@ 5 Discussions
Finally we note that while the asymptotic geometry is taken to be flat space-time (times an internal manifold) the geometries that we can construct in the interior are not limited to flat space-time. For example is ten dimensional type IIB string theory we can take a stack of $N$ spherical D3-branes so that it has finite energy, but take the radius of the sphere to be very large so that the relaxation time is large. In that case the near horizon geometry of a local patch of the D3-branes will be $AdS_{5}\times S^{5}$ with five form flux[9] and we can study properties of this background by sending observers close to the horizon. [^6] If we scale the 4-form field strength according to (2.5) then the three-brane charge will grow as $\lambda^{4}$[1] and we shall get a locally flat space-time. To get $AdS_{5}\times S^{5}$ geometry we need to keep the three brane charge fixed. Of course if we try to approach infinitely close to the horizon any experiment will take infinite amount of time from the asymptotic observer’s point of view due to the red shift, but once we fix the accuracy with which we want to do the experiment and estimate the time that will be needed from the asymptotic observer’s point of view to perform such an experiment, we can take the size of the 3-brane system to be sufficiently large so that it does not evolve appreciably during the period of the experiment. Similarly by taking large stacks of spherical M2-branes or M5-branes in M-theory we can produce $AdS_{4}\times S^{7}$ or $AdS_{7}\times S^{4}$ geometries[9]. Since both, the ten dimensional type IIB string theory and the eleven dimensional M-theory, can be produced as backgrounds in the interior of an asymptotically flat four dimensional string theory, obtained by compactifying type IIA string theory on $T^{6}$, we see that all the $AdS_{p}\times S^{q}$ backgrounds listed above can be considered as different states of the same underlying theory.